what is the total space required to accomodate 10-8 AWG conductors in an electrical box?
A-355 ml
B-422.6 ml
C-444.2 ml
D-451 ml
Answer #1

Option B /422.6 ml (closest)

For 8 to 10 AWG conductors we have 1 conductor volume = 2.58 cubic inch

For 10 conductors = 10*2.58= 25.8cubic inch

25.8 cubic inch =( 25.8 * 16.4) ml = 423 ml